OpenStudy (anonymous):

Two numbers, 4 and a, have a geometric mean of 12. Find the value of a.

OpenStudy (anonymous):

@AcademicGurusInc

OpenStudy (anonymous):

i will give medal

OpenStudy (anonymous):

\[ \sqrt{4*a}=12 \] squaring both sides \[(\sqrt{4*a})^2=12^2\] \[4*a=12^2\] Find a

OpenStudy (anonymous):

do i multiply 4 and 12

OpenStudy (anonymous):

wait it would be 3

OpenStudy (anonymous):

\[4*a = 12^2\] \[4*a = 144\] Now divide both sides by 4 to find a

OpenStudy (anonymous):

a=36

OpenStudy (anonymous):

is it

OpenStudy (anonymous):

yups

OpenStudy (anonymous):

thnks